OpenStudy (anonymous):

Find an exact value cos(19pi/12)

OpenStudy (campbell_st):

ok... this is a 2nd quadrant problem so cos is negative and you need the sum of 2 angles formula \[\cos(\frac{19\pi}{12}) = \cos(\frac{5\pi}{6} + \frac{3\pi}{4})\] so you need to find the 2 exact value triangles \[\cos(\frac{5\pi}{6}) = - \cos(\frac{\pi}{6})\] \[\cos(\frac{3\pi}{4}) = - \cos(\frac{\pi}{4})\] you will also need to find \[\sin(\frac{5\pi}{6}) = \sin(\frac{\pi}{6})....and....\sin(\frac{3\pi}{4}) = \sin(\frac{\pi}{4})\] hope this helps

OpenStudy (anonymous):

is the answer sqrt 6 - sqrt 2 all over 4?

OpenStudy (campbell_st):

I haven't done the calculation

OpenStudy (campbell_st):

yep... thats the answer I got

OpenStudy (anonymous):

okay thank you!